A school yard sandbox is in the shape of a rectangle. It has 54 feet of boards around the border that enclose the sand box. If the length of the sandbox is 3 feet less than twice the width, find the dimensions of the sand box.
Answer by checkley77(12844)

L=2W-3
P=2L+2W
54=2(2W-3)+2W
54=4W-6+2W
6W=54+6
6W=60
W=60/6
W=10 ANS. FOR THE WIDTH.
L=2*10-3
L=20-3
L=17 ANS. FOR THE LENGTH.
PROOF:
54=2*17+2*10
54=34+20
54=54